# Transport alpha divergences

### v2

@@ Abstract.
We derive a class of divergences measuring the difference between probability density functions on a one-dimensional sample space. This divergence is a one-parameter variation of the Itakura–Saito divergence between quantile density functions. We prove that the proposed divergence is one-parameter variation of transport Kullback-Leibler divergence and Hessian distance of negative Boltzmann entropy with respect to Wasserstein-$2$ metric. From Taylor expansions, we also formulate the $3$-symmetric tensor in Wasserstein space, which is given by an iterative Gamma three operators. The alpha-geodesic on Wasserstein space is also derived. From these properties, we name the proposed information measures transport alpha divergences. We provide several examples of transport alpha divergences for generative models in machine learning applications.
@@ 1. Introduction
Information measures of differences between probability distributions play essential roles in statistics, information theory, signal processing, and estimation [1, 7, 10, 11]. It is a generalization of Kullback-Leibler (KL) divergence, with dualities and variational properties. One typical example is the Alpha divergence, which has vast applications in machine learning inference problems and Bayesian sampling problems.
@@
Information geometry (IG) studies the geometric, duality, and invariance properties of divergences in probability density space. Examples include Kullback-Leibler (KL) divergence, alpha-divergences, and their generalizations [1, 7, 10, 31]. In this study, the derivatives of negative Boltzmann–Shannon entropy in $L^{2}$ space are with fundamental roles. Its first-order derivative is the likelihood function, its second derivative forms the Fisher-Rao metric, while its third derivative introduces the Amari–Chenstov tensor. The combination of these derivatives characterizes the KL divergence and its one-parameter variation, namely alpha divergences, where alpha is a scalar. From these characterizations, IG studies and constructs finite-dimensional probability models, with desirable approximation and convexity properties in inference problems.
@@
Recently, optimal transport studies the other distance function, namely Wasserstein distances, in probability density space [30]. The Wasserstein distance introduces the dualities based on pushforward mapping functions, which nowadays have vast applications in estimation and AI sampling problems, such as generative adversarial networks [4]. In particular, there also exists a Riemannian metric for Wasserstein-$2$ distance in probability density space [14, 28, 30]. Under this Wasserstein-$2$ metric, the derivatives of Boltzman-Shannon entropy are of importance in simulating physics equations [23, 13] and Ricci curvature lower bound in a sample space [8, 22, 25]. The study of first- and second-order derivatives in Wasserstein-$2$ space has also been used in statistics and optimization in machine learning algorithms [15]. A natural question arises. What is the alpha divergence in terms of pushforward mapping functions?
@@
This paper answers this question by applying information geometry methods to optimal transport geometry. For simplicity of presentation, we focus on the result in one-dimensional sample space. We show that the transport alpha divergence is a one-parameter family, which interpolates the transport KL divergence function [17] and transport Hessian distances [18]. We derive the third order derivative, a.k.a. 33–symmetric tensor, of negative Boltzmann–Shannon entropy in Wasserstein-$2$ space. Several properties of transport alpha divergences are presented, including duality relation, Taylor expansions, generalized Bregman divergences, and generalized Pythagorean theorem in Wasserstein-$2$ spaces.
@@
We briefly present the main result. Given a one-dimensional domain $\Omega$ and two strictly positive probability density functions $p$, $q$, we propose the alpha divergence in Wasserstein-$2$ space as
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{(}\big{(}\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\big{)}^{\alpha}-{\alpha}\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-1\Big{)}dx,&\qquad\textrm{if $\alpha\neq 0$;}\\ &\frac{1}{2}\int_{0}^{1}|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}|^{2}q(x)dx,&\qquad\textrm{if $\alpha=0$.}\end{aligned}\right$. (1)
@@
where $Q_{p}$, $Q_{q}$ are quantile functions of $p$, $q$, respectively, and $Q^{\prime}_{p}$, $Q^{\prime}_{q}$ are derivatives of quantile functions, namely quantile density functions. We note that the quantile function is the inverse function of cumulative distribution function. We remark that compared with alpha divergences in $L^{2}$ space, the transport alpha divergence reformulates density functions to quantile density functions.
@@
In literature, several joint studies exist among information geometry, optimal transport, and alpha divergences [5, 21, 24, 27]. For example, [24, 27] studies the optimal transport over a Bregman divergence ground cost. [21] studies the matrix decomposition viewpoint for different information metrics on Gaussian families. Recently, [5] also studies Wasserstein-2 metric with general Riemannian ground metric, and then study canniocal divergences from Wasserstein-2 distances. Compared to the above studies, we focus on the Hessian metric of negative Boltzman-Shannon entropies in Wasserstein-$2$ space. In this paper, we apply Hessian structure [1, 26] to construct divergence functionals in terms of Jacobi operators of mapping functions; see related works in [15, 16, 17, 18].
@@
This paper is organized as follows. In section 2, we briefly review the definition of classical alpha divergence in positive octant and its relation with information geometry methods. In section 4, we first construct transport alpha divergence in one-dimensional sample space. Its Taylor expansions show both the Hessian metric and the 3-symmetric tensor in Wasserstein-$2$ space. Properties of transport alpha divergences, including generalized Bregman divergences and Pythagorean theorem in Wasserstein-$2$ spaces, are discussed. Several analytical formulas in generative models are provided in section 5.
@@ 2. Divergence functions and Information geometry methods
In this section, we briefly review alpha divergence functions in positive octant. We also recall information geometry methods for studying these divergences functionals [1, 2, 7].
From now on, we mainly follow derivations in [7, section 2.5.2]. Denote a dd–dimensional positive octant by $(0,+\infty)^{d}$. For any vectors $m=(m_{i})_{i=1}^{d}$, $n=(n_{i})_{i=1}^{d}\in(0,+\infty)^{d}$, the alpha divergence is defined by
$\mathrm{D}_{\alpha}[m\|n]=\sum_{i=1}^{d}f_{\alpha}(\frac{m_{i}}{n_{i}})n_{i},$
@@
We note that, in general when $\alpha\neq 0$, alpha divergence is not a distance function, since $\mathrm{D}_{\alpha}(m|n)\neq\mathrm{D}_{\alpha}(n|m)$. The following dual relation holds
@@
The alpha divergence has several important properties from Hessian structures of an entropy function, especially Taylor expansions and alpha geodesics. Denote a finite dimensional Boltzman-Shannon entropy function by $H(m)=-\sum_{i=1}^{n}m_{i}\log m_{i}$. Denote the Hessian matrix of negative $H$, also named Fisher matrix, by
@@
and denote the third derivative of $H$ by a 3–symmetric tensor, also known as Amari-Chentsov tensor,
@@
where $\delta_{ij}$ is a Kronecker delta function. The above Hessian matrix and $3$-tensor are useful in constructing alpha divergences.
Firstly, the Taylor expansion of alpha divergence holds:
@@
where $|\cdot|$ is the Euclidean norm in $\mathbb{R}^{d}$. Secondly, there are a pair of dual geodesics, namely $\pm\alpha$-geodesics. Denote the $\alpha$-connections at a point $m\in\mathbb{R}_{+}^{d}$ by a three index symbol
@@
Then the alpha geodesic is given below. Denote $\gamma_{\alpha}(t)\in\mathbb{R}_{+}^{d}$, $t\in[0,1]$, with initial point and terminal point $\gamma_{\alpha}(0)=m$, $\gamma_{\alpha}(1)=n$, and
@@
In above formula, the power and product are componentwise. If $\alpha=-1$, then ODE (2)’s solution is named the mixture (m)-geodesics:
@@
With above defined alpha geodesics, there are duality properties of alpha divergences, including Bregman divergences in terms of $\alpha$-representations (3), and generalized Pythagorean theorem. In literature [2, 26, 31], $(\mathbb{R}_{+}^{d},g,T)$ is an example of $\alpha$-geometry, or Hessian structure of entropy function $H$.
@@ 3. Transport alpha divergences
In this section, we define alpha divergences in Wasserstein-$2$ space. Several properties are presented, including composite Bregman divergences and generalized Pythagorean theorem in Wasserstein-$2$ spaces. We also define the alpha geodesic for the completeness of the result.
@@ 3.1. Review of Wasserstein-$2$ distances
We briefly review some basic facts on Wasserstein-$2$ distances [3]. We only consider the one dimensional case. In the statistics literatue, the Wasserstein distance will be represented by the quantile functions [12].
@@
where $\int$, $dx$ are standard integration symbols in $1D$. For any two continuous probability densities $p$, $q\in\mathcal{P}(\Omega)$ with finite second moments, the Wasserstein-$2$ distance [3, 30] is defined by:
@@
where the infimum is taken over all continuous mapping function $T$ that pushforwards $q$ to $p$. In other words, $T_{\#}q=p$, which means the Monge-Amperé equation holds:
@@
In one-dimensional space, the optimal mapping function $T$ is monotone, which can be solved analytically in terms of quantile functions. From now on, we denote the cumulative distribution functions (CDFs) $F_{p}$, $F_{q}$ of probability density function $p$, $q$, such that
@@
Note that $F_{p}$ and $F_{q}$ are strictly monotonic increasing functions. We write $F_{p}^{-1}$, $F_{q}^{-1}$ are inverse CDFs of $p$, $q$, respectively. We are ready to solve equation (5). Taking the integration on both sides of equation (5) w.r.t. $x$, we have
@@
From the inverse function of a CDF, the optimal transport mapping function satisfies
@@
From now on, we always use $T(x)$ to represent the optimal mapping function. Equivalently, the squared Wasserstein-$2$ distance can be formulated as follows.
$\begin{split}\textrm{W}_{2}(p,q)^{2}=&{\int_{\Omega}|Q_{p}(F_{q}(x))-x|^{2}q(x)dx}\\ =&{\int_{0}^{1}|Q_{p}(u)-Q_{q}(u)|^{2}du},\end{split}$
where we apply the change of variable $u=F_{q}(x)\in[0,1]$ in the second equality. In other words, the Wasserstein-$2$ distance in one dimension is the $L^{2}$ distance in quantile functions. In other words, Wasserstein-$2$ geometry is flat since $\mathcal{P}(\Omega)$ is isometric to a convex subset of $L^{2}([0,1])$.
@@ 3.2. Transport alpha divergences
$f_{\mathrm{T},\alpha}(z)=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}(z^{\alpha}-\alpha\log z-1),&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\frac{1}{2}|\log z|^{2},&\quad\mathrm{if}\quad\alpha=0.\end{aligned}\right$.
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q):=\int_{\Omega}f_{\mathrm{T},\alpha}(T^{\prime}(x))q(x)dx=\int_{\Omega}f_{\mathrm{T},\alpha}(\frac{q(x)}{p(T(x))})q(x)dx,$
where $T$ is the monotone mapping function that pushforwards $q$ to $p$, such that $T_{\#}q=p$. We name $\mathrm{D}_{\mathrm{T},\alpha}$ the transport alpha divergence.
We can represent the transport alpha divergence in terms of quantile density functions (QDFs). Denote the QDFs of probability densities $p$ and $q$ below.
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\int_{0}^{1}f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})du.$ (7)
@@
Denote a variable $u=F_{q}(x)$, $u\in[0,1]$. Thus, by changing $x$ to $u$ in the following integration,
$\begin{split}\int_{\Omega}f_{\mathrm{T},\alpha}(T^{\prime}(x))q(x)dx=&\int_{\Omega}f_{\mathrm{T},\alpha}(\frac{d}{dx}Q_{p}(F_{q}(x)))q(x)dx\\ =&\int_{\Omega}f_{\mathrm{T},\alpha}\Big{(}\frac{\frac{d}{du}Q_{p}(u)|_{u=F_{q}(x)}}{1/\frac{dF_{q}(x)}{dx}}\Big{)}q(x)dx\\ =&\int_{0}^{1}f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})du,\end{split}$
@@
    $\mathrm{D}_{\mathrm{T},0}(p\|q)=\frac{1}{2}\int_{0}^{1}|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}|^{2}du.$

### v3

@@ Abstract.
We derive a class of divergences measuring the difference between probability density functions on the one-dimensional sample space. This divergence is a one-parameter variation of the Itakura–Saito divergence between quantile density functions. We prove that the proposed divergence is a one-parameter variation of the transport Kullback-Leibler divergence and the Hessian distance of negative Boltzmann entropy with respect to the Wasserstein-$2$ metric. From Taylor expansions, we also formulate the $3$-symmetric tensor in Wasserstein-$2$ space, which is given by an iterative Gamma three operator. The alpha-geodesic on Wasserstein space is also derived. From these properties, we name the proposed divergences transport alpha divergences. We provide several examples of transport alpha divergences on one dimensional distributions, such as generative models and Cauchy distributions.
@@ 1. Introduction
Divergences between probability distributions play essential roles in statistics, information theory, and signal processing [1, 7, 10, 11]. It is a generalization of Kullback-Leibler (KL) divergence, with dualities and variational properties. One typical example is the alpha divergence, which has vast applications in machine learning inference problems and Bayesian sampling problems.
@@
Information geometry (IG) studies Legendre dualities, and invariance properties of divergences in probability density space. Examples include Kullback–Leibler (KL) divergence, alpha divergences, and their generalizations [1, 7, 10, 31]. In this study, the derivatives of negative Boltzmann–Shannon entropy in $L^{2}$ space play fundamental roles. Its first-order derivative is the likelihood function, its second derivative satisfies the Fisher-Rao metric, while its third derivative introduces the Amari–Chenstov tensor. These derivatives characterize the KL divergence and its one-parameter variation, namely alpha divergences, where alpha is a scalar. From these characterizations, IG studies and constructs finite-dimensional probability models, with approximation and convexity properties in inference problems.
@@
Recently, optimal transport leads to a different kind of divergences between probability distributions, of which the Wasserstein distance is a fundamental example [30]. The Wasserstein distance introduces the dualities based on the ground cost function, known as the Kantorovich duality. Optimal transport nowadays has vast applications in artificial intelligence (AI), such as generative adversarial networks [4]. In particular, there also exists a Riemannian metric for Wasserstein-$2$ distance in probability density space [15, 28, 30]. Under the Wasserstein-$2$ metric, the derivatives of negative Boltzman-Shannon entropy are of importance in simulating physics equations [14] and Ricci curvature lower bound in a sample space [8, 13, 23, 25]. The study of first- and second-order derivatives in Wasserstein-$2$ space has also been used in statistics and optimization with applications in machine learning algorithms [16]. A natural question arises. Is there an analogue of the alpha divergence based on optimal transport?
@@
This paper partially addresses this question. We apply information geometry methods, such as the dualistic geometry of a divergence/contrast function, to construct divergences of probability densities from transport maps. For simplicity of presentation, we focus on the result in one-dimensional sample space, where the ground distance is the squared Euclidean distance. We introduce a one-parameter family, named transport alpha divergence, which interpolates the transport KL divergence [18] and the transport Hessian distance [19]. We derive the third order derivative, i.e., a 33–symmetric tensor, of the negative Boltzmann–Shannon entropy in Wasserstein-$2$ space. Several properties of transport alpha divergences are presented, including dualities, Taylor expansions, generalized Bregman divergences, and the generalized Pythagorean theorem in Wasserstein-$2$ space.
@@
We briefly present the main result. Given a one-dimensional domain $\Omega$ and two strictly-positive continuous probability density functions $p$, $q$, we define the transport alpha divergence as
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{0}^{1}\Big(\big(\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\big)^{\alpha}-{\alpha}\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-1\Big){du},&\qquad\textrm{if $\alpha\neq 0$;}\\ &\frac{1}{2}\int_{0}^{1}\left|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\right|^{2}{du},&\qquad\textrm{if $\alpha=0$.}\end{aligned}\right$. (1)
@@
where $Q_{p}$, $Q_{q}$ are quantile functions of densities $p$, $q$, respectively, and $Q^{\prime}_{p}$, $Q^{\prime}_{q}$ are derivatives of quantile functions, namely quantile density functions. We note that the quantile function is the inverse function of cumulative distribution function. We remark that compared with alpha divergences in $L^{2}$ space, the transport alpha divergence studies the difference between quantile density functions, instead of probability density functions.
@@
In literature, several joint studies exist among information geometry, optimal transport, and alpha divergences [5, 22, 24, 27]. For example, [27] uses OT duality to generalize the Bregman divergence. [22] studies the matrix decomposition viewpoint for different information metrics on Gaussian families. Recently, [5] also studies Wasserstein-2 metric with general Riemannian ground metric, and then study canniocal divergences from Wasserstein-2 distances. Compared to the above studies, we focus on the Hessian metric of the negative Boltzman-Shannon entropy in Wasserstein-$2$ space. In this paper, we apply Hessian structures [1, 26] to construct divergence functionals in terms of derivatives of transport maps; see related works in [16, 17, 18, 19].
@@
This paper is organized as follows. In section 2, we briefly review the definition of classical alpha divergence in positive octant and its relation with information geometry methods. In section 4, we first construct transport alpha divergence in one-dimensional sample space, whose Taylor expansion shows both the Hessian metric and the 3-symmetric tensor in Wasserstein-$2$ space. Properties of transport alpha divergences, including generalized Bregman divergences and Pythagorean theorem in Wasserstein-$2$ space, are discussed. Several analytical formulas in generative models are provided in section 5.
@@ 2. Divergence functions and Information geometry methods
In this section, we briefly review alpha divergences in positive octant, and present information geometry methods for studying these divergences [1, 2, 7]. E.g., we mainly follow [7, section 2.5.2] for the definition of alpha connection.
Denote a dd–dimensional positive octant by $(0,+\infty)^{d}$. For any vectors $m=(m_{i})_{i=1}^{d}$, $n=(n_{i})_{i=1}^{d}\in(0,+\infty)^{d}$, the alpha divergence is defined by
$\mathrm{D}_{\alpha}[m\|n]=\sum_{i=1}^{d}f_{\alpha}\left(\frac{m_{i}}{n_{i}}\right)n_{i},$
@@
We note that, in general when $\alpha\neq 0$, alpha divergence is not a distance function, since $\mathrm{D}_{\alpha}[m\|n]\neq\mathrm{D}_{\alpha}[n\|m]$. The following dual relation holds
@@
The alpha divergence has several important properties from the Hessian structure of an entropy function, including Taylor expansions and alpha geodesics. Denote a finite dimensional Boltzman-Shannon entropy function by $H(m)=-\sum_{i=1}^{n}m_{i}\log m_{i}$. Denote the Hessian matrix of negative $H$, also named Fisher matrix, by
@@
and denote the third derivative of $H$ by a 3–symmetric tensor, known as Amari-Chentsov tensor,
@@
where $\delta_{ij}$ is a Kronecker delta function. The above Hessian matrix and $3$-tensor are useful in studying the alpha divergences.
Firstly, the Taylor expansion of alpha divergences hold:
@@
[…]
@@
Then the alpha geodesic is given below. Denote $\gamma_{\alpha}(t)\in\mathbb{R}_{+}^{d}$, $t\in[0,1]$, with both initial and terminal points $\gamma_{\alpha}(0)=m$, $\gamma_{\alpha}(1)=n$, and
@@
In above formula, the power and product are componentwise. In other words, $\gamma_{\alpha}(t)_{k}=\Big((1-t)m_{k}^{\frac{1-\alpha}{2}}+tn_{k}^{\frac{1-\alpha}{2}}\Big)^{\frac{2}{1-\alpha}}$. If $\alpha=-1$, then ODE (2)’s solution is named the mixture (m)-geodesics:
@@
With above defined alpha geodesics, there are duality properties of alpha divergences, including Bregman divergences in terms of $\alpha$-representations (3), and generalized Pythagorean theorem. In literature [2, 7, 26, 31], $(\mathbb{R}_{+}^{d},g,T)$ is an example of $\alpha$-geometry, or Hessian structure of entropy function $H$.
Two remarks are presented here. The $\alpha$-geometry in positive probability measures are in general different from the alpha geometry in positive measures, where there is a projection to be studied on positive probability measures. In addition, the information geometry method is not limit to the entropy function. See details in [7].
@@ 3. Transport alpha divergences
In this section, we define alpha divergences in one dimensional Wasserstein-$2$ space. Several properties are presented, including composite Bregman divergences and generalized Pythagorean theorem in Wasserstein-$2$ space. We also define the alpha geodesic for the completeness of the result.
@@ 3.1. Review of Wasserstein-$2$ distances
We briefly review some basic facts on Wasserstein-$2$ distances. We only consider the one dimensional case. In this case, the Wasserstein distance will be represented by quantile functions [12].
@@
where $\int$, $dx$ are standard integration symbols in $1D$. For any two continuous probability densities $p$, $q\in\mathcal{P}(\Omega)$ with finite second moments, the Wasserstein-$2$ distance [30] is defined in the Monge problem by:
@@
where the infimum is taken over the continuous transport map function $T$ that pushforwards $q$ to $p$. When transport map function $T$ is differentiable, $T_{\#}q=p$ means that the Monge-Amperé equation holds:
@@
In one-dimensional space, the optimal transport map function $T$ is increasing, which can be solved analytically in terms of quantile functions. From now on, we denote the cumulative distribution functions (CDFs) $F_{p}$, $F_{q}$ of probability density function $p$, $q$, such that
@@
Note that $F_{p}$ and $F_{q}$ are strictly increasing functions. We write $F_{p}^{-1}$, $F_{q}^{-1}$ are inverse CDFs of $p$, $q$, respectively. We are ready to solve equation (5). Taking the integration on both sides of equation (5) w.r.t. $x$, we have
@@
Thus, the optimal transport map function satisfies
@@
From now on, we always use $T(x)$ to represent the optimal transport map. Equivalently, the squared Wasserstein-$2$ distance can be formulated as follows.
$\begin{split}\textrm{W}_{2}(p,q)^{2}=&{\int_{\Omega}|Q_{p}(F_{q}(x))-x|^{2}q(x)dx}={\int_{0}^{1}|Q_{p}(u)-Q_{q}(u)|^{2}du},\end{split}$
where we apply the change of variable $u=F_{q}(x)\in[0,1]$ in the second equality. In other words, the Wasserstein-$2$ distance in one dimension is the $L^{2}$ distance in quantile functions. In other words, the one-dimensional Wasserstein-$2$ geometry is flat since $(\mathcal{P}(\Omega),W_{2})$ is isometric to a convex subset of $L^{2}([0,1])$.
@@ 3.2. Transport alpha divergences
$f_{\mathrm{T},\alpha}(z)=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\left(z^{\alpha}-\alpha\log z-1\right),&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\frac{1}{2}|\log z|^{2},&\quad\mathrm{if}\quad\alpha=0.\end{aligned}\right$.
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q):=\int_{\Omega}f_{\mathrm{T},\alpha}(T^{\prime}(x))q(x)dx=\int_{\Omega}f_{\mathrm{T},\alpha}\left(\frac{q(x)}{p(T(x))}\right)q(x)dx,$
where $T$ is the monotone transport map function that pushforwards $q$ to $p$, such that $T_{\#}q=p$. We name $\mathrm{D}_{\mathrm{T},\alpha}$ the transport alpha divergence.
From now on, we assume that $f_{\mathrm{T},\alpha}(T^{\prime}(x))$ is integrable with the weight function $q(x)$. This ensures that $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)<+\infty$. We can represent the transport alpha divergence in terms of quantile density functions (QDFs). Denote the QDFs of probability densities $p$ and $q$ below.
@@
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\int_{0}^{1}f_{\mathrm{T},\alpha}\left(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\right)du.$ (7)
@@
Denote a variable $u=F_{q}(x)$, $u\in[0,1]$. Thus, by changing $x$ to $u$ in the following integration, we have
$\begin{split}\int_{\Omega}f_{\mathrm{T},\alpha}(T^{\prime}(x))q(x)dx=&\int_{\Omega}f_{\mathrm{T},\alpha}\left(\frac{d}{dx}Q_{p}(F_{q}(x))\right)q(x)dx\\ =&\int_{\Omega}f_{\mathrm{T},\alpha}\Big(\frac{\frac{d}{du}Q_{p}(u)|_{u=F_{q}(x)}}{1/\frac{dF_{q}(x)}{dx}}\Big)q(x)dx\\ =&\int_{0}^{1}f_{\mathrm{T},\alpha}\left(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\right)du,\end{split}$
@@
Remark 1.
We provide a sufficient condition that $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)$ is finite. Assume that there exists constants $C_{p,q}>c_{p,q}>0$, such that $\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\in[c_{p,q},C_{p,q}]$, for almost everywhere $u\in[0,1]$. Then $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)<+\infty$. One example of this condition is the Cauchy distribution; see Example 3.
@@
    $\mathrm{D}_{\mathrm{T},0}(p\|q)=\frac{1}{2}\int_{0}^{1}\left|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\right|^{2}du.$